# Low-Luminosity Type IIP Supernovae from the Zwicky Transient Facility Census of the Local Universe. II: Lightcurve Analysis

### v1

@@ 7.5 What are the progenitors of LLIIP SNe?
7.5.2 sAGB stars that undergo Electron-capture SNe
Progenitors in the mass range may evolve into super-Asymptotic Giant Branch (sAGB) stars and explode as electron-capture supernovae (ECSNe). ECSN simulations consistently produce low explosion energies and nickel yields significantly lower than typical Fe CCSNe but similar to those observed for the lowest-mass Fe CCSNe. According to Sato et al. (2024), ECSNe, without significant circumstellar material (CSM) interaction, exhibit distinctively bluer bolometric lightcurve plateaus compared to Fe CCSNe, characterized by mid-plateau colors exceeding the relation . This difference arises from the lower-density envelopes inherent to sAGB progenitors. We plot the mid-plateau colors for our SN sample against the Sato et al. (2024) criterion in Figure 18. While SNe 2022omr, 2021tyw, 2021cwe, and 2022prv show bluer mid-plateau colors, their peak luminosities and nickel masses favor an Fe CCSN origin. Among our LLIIP SNe sample, SNe 2023wcr, 2021gmj, 2021zgm, and 2020abcq exhibit notably bluer mid-plateau colors. However, we note that the mid-plateau colors might also be affected by significant CSM interaction. The lightcurves of the SNe with the bluest (SN 2021tyw, mid-plateau mag) and reddest (SN 2022jzc, mid-plateau mag) colors are shown in Figure 18.
@@ 8 Conclusion
9 Acknowldegement
@@
ZTF22abkhrkd | SN 2022wol | 10.24 0.43 | |||||
ZTF22abtjefa | SN 2022aaad | 11.03 0.63 | |||||
ZTF19abqrhvt | SN 2019nyk | 11.69 0.86 | |||||
ZTF19acewuwn | SN 2019ssl | 12.56 1.09 | |||||
ZTF21aaqugxm | SN 2021hdt | 14.82 1.01 | |||||
ZTF24aaasazz | SN 2024ov | 10.47 0.50 | |||||
ZTF20abyylgi | SN 2020svn | 13.38 1.42 | |||||
ZTF24aagupsf | SN 2024egd | 12.03 0.51 | |||||
ZTF21aanzcuj | SN 2021enz | 13.08 0.33 | |||||
ZTF20acveadu | SN 2020absk | 10.31 0.33 | |||||
ZTF19aamwhat | SN 2019bzd | 10.51 0.66 | |||||
ZTF19abbwfgp | SN 2019ikb | 12.37 1.15 | |||||
ZTF23aanymcl | SN 2023kzz | 11.63 0.73 | |||||
ZTF18aawpwlf | SN 2020hvn | 10.64 0.68 | |||||
ZTF18aatyqds | SN 2018btl | 12.31 0.45 | |||||
ZTF18aaxkqgy | SN 2018ccb | 12.06 1.13 | |||||
ZTF21acafqtj | SN 2021yok | 12.63 0.93 | |||||
ZTF20aaiovoy | SN 2020bad | 10.45 3.45 | |||||
ZTF21aafepon | SN 2021ass | 11.00 0.57 | |||||
ZTF23abhzfww | SN 2023twg | 11.96 0.64 | |||||
ZTF20abyzomt | SN 2020szs | 12.08 0.62 | |||||
ZTF24aaucrua | SN 2024nez | 9.86 0.49 | |||||
ZTF21aaobkmg | SN 2021eui | 10.54 0.48 | |||||
ZTF21aagtqna | SN 2021brb | 13.02 0.81 | |||||
ZTF19aadnxnl | SN 2019va | 14.05 0.25 | |||||
ZTF19aanrrqu | SN 2019clp | 12.11 0.94 | |||||
ZTF20aciojzg | AT 2020wip | 10.43 0.50 | |||||
ZTF22abssiet | SN 2022zmb | 10.53 0.88 | |||||
ZTF18aaqfqaa | SN 2018bek | 10.00 0.55 | |||||
ZTF19aailepg | SN 2019amt | 12.55 0.48 | |||||
ZTF23abaxtlq | SN 2023rix | 10.70 0.34 | |||||
ZTF21aapkcmr | AT 2021fnj | 11.94 1.00 | |||||
ZTF18aawgrxz | SN 2021lmp | 10.81 0.89 | |||||
ZTF21acgrrnl | SN 2021aayf | 13.89 2.66 | |||||
ZTF22abkbjsb | SN 2022vym | 10.95 0.65 | |||||
ZTF19aclobbu | SN 2019twk | 11.55 0.97 | |||||
ZTF22aafsqud | SN 2022hql | 11.93 0.89 | |||||
ZTF22abfavpu | SN 2022tmb | 14.35 2.06 | |||||
ZTF21abfiuqf | SN 2021pla | 10.50 1.47 | |||||
ZTF24aarvbxj | SN 2024lby | 11.73 1.00 | |||||
ZTF24aaemydm | SN 2024chx | 10.17 0.59 | |||||
ZTF22abtspsw | SN 2022aagp | 11.40 1.62 | |||||
ZTF20acfvprb | AT 2020thp | 12.12 0.49 | |||||
ZTF23aasbvab | SN 2023ngy | 11.80 0.76 | |||||
ZTF22aaevwec | SN 2022gwg | 11.03 1.23 | |||||
ZTF21acfajbc | SN 2021aalq | 10.77 0.66 | |||||
ZTF24aadkwni | SN 2024aul | 10.32 1.02 | |||||
ZTF21abouuat | SN 2021ucg | 16.65 1.01 |
@@
ZTF18aahfqbc | SN 2018aoq | 12.05 0.78 | |||||
ZTF21abhhrpj | SN 2021qiu | 15.71 2.99 | |||||
ZTF20actnuls | SN 2020aawo | 12.06 0.57 | |||||
ZTF22abnujbv | SN 2022xus | 10.57 1.34 | |||||
ZTF23abbtkrv | SN 2023rvo | 15.79 0.58 | |||||
ZTF19aaniore | SN 2019ceg | 12.19 1.24 | |||||
ZTF22aasojye | SN 2022omr | 13.16 0.51 | |||||
ZTF20aafckit | SN 2020uu | 10.24 0.99 | |||||
ZTF21aaagypx | SN 2021V | 14.23 1.12 | |||||
ZTF19acbwejj | SN 2019upq | 14.86 2.75 | |||||
ZTF23aackjhs | SN 1995al | 9.46 0.70 | |||||
ZTF21aaeqwov | AT 2021htp | 10.40 1.14 | |||||
ZTF19aanhhal | SN 2019cec | 14.22 3.80 | |||||
ZTF22aanrqje | SN 2022mji | 9.41 0.67 | |||||
ZTF21aafkwtk | SN 2021apg | 12.42 0.36 | |||||
ZTF24abmkros | SN 2024xkd | 10.82 1.04 | |||||
ZTF22aaolwsd | SN 2022mxv | 11.75 0.81 | |||||
ZTF24abtczty | SN 2024abfl | 10.71 0.45 | |||||
ZTF19abajxet | SN 2019hyk | 11.93 0.55 | |||||
ZTF24aaplfjd | SN 2024jxm | 10.39 0.34 | |||||
ZTF21aantsla | SN 2021ech | 11.06 0.57 | |||||
ZTF24aajxppf | SN 2024grw | 10.96 0.93 | |||||
ZTF19acftfav | SN 2019ssi | 10.47 0.37 | |||||
ZTF20actkmks | SN 2020aaun | 9.98 2.83 | |||||
ZTF21abnlhxs | SN 2021tyw | 14.43 1.29 | |||||
ZTF23aaphnyz | SN 2023lkw | 12.80 0.32 | |||||
ZTF19actnwtn | SN 2019vdl | 12.05 0.48 | |||||
ZTF18aadsuxd | SN 2020yui | 12.03 1.36 | |||||
ZTF22aativsd | SN 2022ovb | 12.08 0.40 | |||||
ZTF23aaqknaw | SN 2023lzn | 13.69 0.73 | |||||
ZTF24aabsmvc | SN 2024ws | 11.32 0.87 | |||||
ZTF22abvaetz | SN 2022aang | 10.21 0.66 | |||||
ZTF22ablnrcv | SN 2022xav | 9.28 0.68 | |||||
ZTF21aardvtn | AT 2021htp | 10.22 0.23 | |||||
ZTF22aauurbv | SN 2022pgf | 10.37 0.36 | |||||
ZTF19aazyvub | SN 2019hnl | 11.61 0.51 | |||||
ZTF24aaejecr | SN 2024btj | 10.54 0.83 | |||||
ZTF22aakdqqg | SN 2022kad | 11.95 0.76 | |||||
ZTF22aarycqo | SN 2022ojo | 13.18 1.81 | |||||
ZTF23abascqa | SN 2023rbk | 10.28 1.02 | |||||
ZTF23aasrcyv | SN 2023nlu | 11.00 1.15 | |||||
ZTF23aailjjs | SN 2023hcp | 11.81 0.46 | |||||
ZTF21aabygea | SN 2021os | 12.50 1.34 | |||||
ZTF20acvjagm | SN 2020abtf | 12.73 0.36 | |||||
ZTF19actnyae | SN 2019vdm | 10.50 0.99 | |||||
ZTF21ablvzhp | SN 2021tiq | 10.09 0.68 | |||||
ZTF21aavhnpk | SN 2021jsf | 9.59 1.34 | |||||
ZTF22aaudjgc | SN 2022pfx | 10.45 0.71 |
@@
ZTF22abfxkdm | SN 2022ubb | 10.62 0.43 | |||||
ZTF22aavbfhz | AT 2022phi | 10.64 1.08 | |||||
ZTF22aakdbia | SN 2022jzc | 9.92 0.15 | |||||
ZTF22aazmrpx | SN 2022raj | 10.94 0.50 | |||||
ZTF20acfkzcg | AT 2021htp | 11.40 0.77 | |||||
ZTF21abrluay | SN 2021vfh | 12.92 1.37 | |||||
ZTF19acytcsg | SN 2019wvz | 15.04 0.64 | |||||
ZTF19abwztsb | SN 2019pjs | 10.72 0.76 | |||||
ZTF22aaywnyg | SN 2022pru | 10.86 0.89 | |||||
ZTF19abbnamr | SN 2019iex | 14.16 0.73 | |||||
ZTF23abbsxzs | SN 2023rtq | 9.84 1.38 | |||||
ZTF22abzqwmp | SN 2022adth | 11.95 0.56 | |||||
ZTF20acitoie | SN 2020wgp | 10.74 0.97 | |||||
ZTF19aapafit | SN 2019cvz | 14.95 1.07 | |||||
ZTF20acbjgoa | SN 2020acac | 10.48 0.64 | |||||
ZTF21abvcxel | SN 2021wvw | 9.56 0.57 | |||||
ZTF20aapchqy | SN 2020cxd | 9.90 0.08 | |||||
ZTF18aaszvfn | SN 2021iaw | 12.27 0.62 | |||||
ZTF20acuhren | SN 2020abcq | 10.28 0.97 | |||||
ZTF23abnogui | SN 2023wcr | 10.75 1.02 | |||||
ZTF22abyivoq | SN 2022acko | 10.97 0.25 | |||||
ZTF22abhsxph | SN 2022vyc | 12.20 0.91 |

### v2

@@ VII.5 What are the progenitors of LLIIP SNe?
VII.5.2 sAGB stars that undergo Electron-capture SNe
Progenitors in the mass range may evolve into super-Asymptotic Giant Branch (sAGB) stars and explode as electron-capture supernovae (ECSNe). ECSN simulations consistently produce low explosion energies and nickel yields significantly lower than typical Fe CCSNe but similar to those observed for the lowest-mass Fe CCSNe. According to Sato et al. (2024), ECSNe, without significant circumstellar material (CSM) interaction, exhibit distinctively bluer bolometric lightcurve plateaus compared to Fe CCSNe, characterized by mid-plateau colors exceeding the relation . This difference arises from the lower-density envelopes inherent to sAGB progenitors. We plot the mid-plateau colors for our SN sample against the Sato et al. (2024) criterion in Figure 18. While SNe 2021tyw, 2022omr, 2021cwe, and 2022prv show bluer mid-plateau colors, their peak luminosities and nickel masses favor an Fe CCSN origin. Among our LLIIP SNe sample, SNe 2023wcr, 2021gmj, 2021zgm, and 2020abcq exhibit notably bluer mid-plateau colors. However, we note that the mid-plateau colors may also be influenced by CSM interaction, as supported by the interaction signatures seen in the spectrum of SN 2021tyw. The lightcurves of the SNe with the bluest (SN 2021tyw, mid-plateau mag) and reddest (SN 2022jzc, mid-plateau mag) colors are shown in Figure 18.
@@ VIII Conclusion
IX Acknowldegement
We thank the anonymous referee for their constructive feedback, which helped improve the quality of this manuscript.
@@ IX Acknowldegement
| ZTF22abkhrkd | SN 2022wol | 10.2 0.4 | |||||
| ZTF22abtjefa | SN 2022aaad | 11.0 0.6 | |||||
| ZTF19abqrhvt | SN 2019nyk | 11.7 0.9 | |||||
| ZTF19acewuwn | SN 2019ssl | 12.6 1.1 | |||||
| ZTF21aaqugxm | SN 2021hdt | 14.8 1.0 | |||||
| ZTF24aaasazz | SN 2024ov | 10.5 0.5 | |||||
| ZTF20abyylgi | SN 2020svn | 13.4 1.4 | |||||
| ZTF24aagupsf | SN 2024egd | 12.0 0.5 | |||||
| ZTF21aanzcuj | SN 2021enz | 13.1 0.3 | |||||
| ZTF20acveadu | SN 2020absk | 10.3 0.3 | |||||
| ZTF19aamwhat | SN 2019bzd | 10.5 0.7 | |||||
| ZTF19abbwfgp | SN 2019ikb | 12.4 1.2 | |||||
| ZTF23aanymcl | SN 2023kzz | 11.6 0.7 | |||||
| ZTF18aawpwlf | SN 2020hvn | 10.6 0.7 | |||||
| ZTF18aatyqds | SN 2018btl | 12.3 0.5 | |||||
| ZTF18aaxkqgy | SN 2018ccb | 12.1 1.1 | |||||
| ZTF21acafqtj | SN 2021yok | 12.6 0.9 | |||||
| ZTF20aaiovoy | SN 2020bad | 10.5 3.4 | |||||
| ZTF21aafepon | SN 2021ass | 11.0 0.6 | |||||
| ZTF23abhzfww | SN 2023twg | 12.0 0.6 | |||||
| ZTF20abyzomt | SN 2020szs | 12.1 0.6 | |||||
| ZTF24aaucrua | SN 2024nez | 9.9 0.5 | |||||
| ZTF21aaobkmg | SN 2021eui | 10.5 0.5 | |||||
| ZTF21aagtqna | SN 2021brb | 13.0 0.8 | |||||
| ZTF19aadnxnl | SN 2019va | 14.0 0.2 | |||||
| ZTF19aanrrqu | SN 2019clp | 12.1 0.9 | |||||
| ZTF20aciojzg | AT 2020wip | 10.4 0.5 | |||||
| ZTF22abssiet | SN 2022zmb | 10.5 0.9 | |||||
| ZTF18aaqfqaa | SN 2018bek | 10.0 0.5 | |||||
| ZTF19aailepg | SN 2019amt | 12.6 0.5 | |||||
| ZTF23abaxtlq | SN 2023rix | 10.7 0.3 | |||||
| ZTF21aapkcmr | AT 2021fnj | 11.9 1.0 | |||||
| ZTF18aawgrxz | SN 2021lmp | 10.8 0.9 | |||||
| ZTF21acgrrnl | SN 2021aayf | 13.9 2.7 | |||||
| ZTF22abkbjsb | SN 2022vym | 11.0 0.6 | |||||
| ZTF19aclobbu | SN 2019twk | 11.6 1.0 | |||||
| ZTF22aafsqud | SN 2022hql | 11.9 0.9 | |||||
| ZTF22abfavpu | SN 2022tmb | 14.3 2.1 | |||||
| ZTF21abfiuqf | SN 2021pla | 10.5 1.5 | |||||
| ZTF24aarvbxj | SN 2024lby | 11.7 1.0 | |||||
| ZTF24aaemydm | SN 2024chx | 10.2 0.6 | |||||
| ZTF22abtspsw | SN 2022aagp | 11.4 1.6 | |||||
| ZTF20acfvprb | AT 2020thp | 12.1 0.5 | |||||
| ZTF23aasbvab | SN 2023ngy | 11.8 0.8 | |||||
| ZTF22aaevwec | SN 2022gwg | 11.0 1.2 | |||||
| ZTF21acfajbc | SN 2021aalq | 10.8 0.7 | |||||
| ZTF24aadkwni | SN 2024aul | 10.3 1.0 | |||||
| ZTF21abouuat | SN 2021ucg | 16.7 1.0 |
@@
| ZTF18aahfqbc | SN 2018aoq | 12.0 0.8 | |||||
| ZTF21abhhrpj | SN 2021qiu | 15.7 3.0 | |||||
| ZTF20actnuls | SN 2020aawo | 12.1 0.6 | |||||
| ZTF22abnujbv | SN 2022xus | 10.6 1.3 | |||||
| ZTF23abbtkrv | SN 2023rvo | 15.8 0.6 | |||||
| ZTF19aaniore | SN 2019ceg | 12.2 1.2 | |||||
| ZTF22aasojye | SN 2022omr | 13.2 0.5 | |||||
| ZTF20aafckit | SN 2020uu | 10.2 1.0 | |||||
| ZTF21aaagypx | SN 2021V | 14.2 1.1 | |||||
| ZTF19acbwejj | SN 2019upq | 14.9 2.7 | |||||
| ZTF23aackjhs | SN 2023bvj | 9.5 0.7 | |||||
| ZTF21aaeqwov | AT 2021htp | 10.4 1.1 | |||||
| ZTF19aanhhal | SN 2019cec | 14.2 3.8 | |||||
| ZTF22aanrqje | SN 2022mji | 9.4 0.7 | |||||
| ZTF21aafkwtk | SN 2021apg | 12.4 0.4 | |||||
| ZTF24abmkros | SN 2024xkd | 10.8 1.0 | |||||
| ZTF22aaolwsd | SN 2022mxv | 11.7 0.8 | |||||
| ZTF24abtczty | SN 2024abfl | 10.7 0.4 | |||||
| ZTF19abajxet | SN 2019hyk | 11.9 0.5 | |||||
| ZTF24aaplfjd | SN 2024jxm | 10.4 0.3 | |||||
| ZTF21aantsla | SN 2021ech | 11.1 0.6 | |||||
| ZTF24aajxppf | SN 2024grw | 11.0 0.9 | |||||
| ZTF19acftfav | SN 2019ssi | 10.5 0.4 | |||||
| ZTF20actkmks | SN 2020aaun | 10.0 2.8 | |||||
| ZTF21abnlhxs | SN 2021tyw | 14.4 1.3 | |||||
| ZTF23aaphnyz | SN 2023lkw | 12.8 0.3 | |||||
| ZTF19actnwtn | SN 2019vdl | 12.0 0.5 | |||||
| ZTF18aadsuxd | SN 2020yui | 12.0 1.4 | |||||
| ZTF22aativsd | SN 2022ovb | 12.1 0.4 | |||||
| ZTF23aaqknaw | SN 2023lzn | 13.7 0.7 | |||||
| ZTF24aabsmvc | SN 2024ws | 11.3 0.9 | |||||
| ZTF22abvaetz | SN 2022aang | 10.2 0.7 | |||||
| ZTF22ablnrcv | SN 2022xav | 9.3 0.7 | |||||
| ZTF21aardvtn | AT 2021htp | 10.2 0.2 | |||||
| ZTF22aauurbv | SN 2022pgf | 10.4 0.4 | |||||
| ZTF19aazyvub | SN 2019hnl | 11.6 0.5 | |||||
| ZTF24aaejecr | SN 2024btj | 10.5 0.8 | |||||
| ZTF22aakdqqg | SN 2022kad | 12.0 0.8 | |||||
| ZTF22aarycqo | SN 2022ojo | 13.2 1.8 | |||||
| ZTF23abascqa | SN 2023rbk | 10.3 1.0 | |||||
| ZTF23aasrcyv | SN 2023nlu | 11.0 1.2 | |||||
| ZTF23aailjjs | SN 2023hcp | 11.8 0.5 | |||||
| ZTF21aabygea | SN 2021os | 12.5 1.3 | |||||
| ZTF20acvjagm | SN 2020abtf | 12.7 0.4 | |||||
| ZTF19actnyae | SN 2019vdm | 10.5 1.0 | |||||
| ZTF21ablvzhp | SN 2021tiq | 10.1 0.7 | |||||
| ZTF21aavhnpk | SN 2021jsf | 9.6 1.3 | |||||
| ZTF22aaudjgc | SN 2022pfx | 10.4 0.7 |
@@
| ZTF22abfxkdm | SN 2022ubb | 10.6 0.4 | |||||
| ZTF22aavbfhz | AT 2022phi | 10.6 1.1 | |||||
| ZTF22aakdbia | SN 2022jzc | 9.9 0.1 | |||||
| ZTF22aazmrpx | SN 2022raj | 10.9 0.5 | |||||
| ZTF20acfkzcg | AT 2021htp | 11.4 0.8 | |||||
| ZTF21abrluay | SN 2021vfh | 12.9 1.4 | |||||
| ZTF19acytcsg | SN 2019wvz | 15.0 0.6 | |||||
| ZTF19abwztsb | SN 2019pjs | 10.7 0.8 | |||||
| ZTF22aaywnyg | SN 2022pru | 10.9 0.9 | |||||
| ZTF19abbnamr | SN 2019iex | 14.2 0.7 | |||||
| ZTF23abbsxzs | SN 2023rtq | 9.8 1.4 | |||||
| ZTF22abzqwmp | SN 2022adth | 11.9 0.6 | |||||
| ZTF20acitoie | SN 2020wgp | 10.7 1.0 | |||||
| ZTF19aapafit | SN 2019cvz | 14.9 1.1 | |||||
| ZTF20acbjgoa | SN 2020acac | 10.5 0.6 | |||||
| ZTF21abvcxel | SN 2021wvw | 9.6 0.6 | |||||
| ZTF20aapchqy | SN 2020cxd | 9.9 0.1 | |||||
| ZTF18aaszvfn | SN 2021iaw | 12.3 0.6 | |||||
| ZTF20acuhren | SN 2020abcq | 10.3 1.0 | |||||
| ZTF23abnogui | SN 2023wcr | 10.8 1.0 | |||||
| ZTF22abyivoq | SN 2022acko | 11.0 0.2 | |||||
| ZTF22abhsxph | SN 2022vyc | 12.2 0.9 |